OpenStudy (anonymous):

Hey there I need help with this distance btw 2 points.

OpenStudy (anonymous):

\[(\sqrt{3}, \sqrt{2}) and (\frac{ 2 }{ \sqrt{3}} ,-\sqrt{2})\]

OpenStudy (kc_kennylau):

\[d=\sqrt{(y_2-y_1)^2-(x_2-x_1)^2}\]

OpenStudy (anonymous):

I did that and I got two different answers than the book

OpenStudy (anonymous):

I have no clue how to solve this correctly

OpenStudy (kc_kennylau):

What did you get? :)

OpenStudy (anonymous):

waitt. Kenny how come you are subtracting them? you are suppose to add them

OpenStudy (anonymous):

No harsh the answer in not 1

OpenStudy (kc_kennylau):

oh sorry

OpenStudy (kc_kennylau):

\[d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\]

OpenStudy (anonymous):

"...got two different answers than the book"

OpenStudy (anonymous):

i got the ans 11/3

OpenStudy (kc_kennylau):

\[\sqrt{(-\sqrt2-\sqrt2)^2+(\frac2{\sqrt3}-\sqrt3)^2}\]\[=\sqrt{8+\frac13}\]\[=\sqrt{\frac{75}9}\]\[=\frac{5\sqrt3}3\]

OpenStudy (anonymous):

Still wrong Harsh

OpenStudy (anonymous):

Correct answer Kenny but...did you distribute?

OpenStudy (anonymous):

OHH! well Kenny Thank you again for helping =) i need a break...

OpenStudy (kc_kennylau):

\[\sqrt{(-\sqrt2-\sqrt2)^2+(\frac2{\sqrt3}-\sqrt3)^2}\]\[=\sqrt{(-2\sqrt2)^2+(\frac{-1}{\sqrt3})^2}\]\[=\sqrt{8+\frac13}\]

OpenStudy (kc_kennylau):

no problem :)

OpenStudy (anonymous):

Kenny at first glance I thought I understood, now I am confused. What did you do?

OpenStudy (kc_kennylau):

Where did you not understand?

OpenStudy (kc_kennylau):

Sorry I have to sleep now, please bump to the top so that people can help you

OpenStudy (anonymous):

oK good night